# Multimodal Graph Benchmark

### v1

@@ Abstract
Associating unstructured data with structured information is crucial for real-world tasks that require relevance search. However, existing graph learning benchmarks often overlook the rich semantic information associate with each node. To bridge such gap, we introduce the Multimodal Graph Benchmark (MM-Graph), the first comprehensive multi-modal graph benchmark that incorporates both textual and visual information. MM-Graph surpasses previous efforts, which have primarily focused on text-attributed graphs with various connectivity patterns. MM-Graph consists of five graph learning datasets of various scales that are appropriate for different learning tasks. Their multimodal node features, enabling a more comprehensive evaluation of graph learning algorithms in real-world scenarios. To facilitate research on multimodal graph learning, we further provide an extensive study on the performance of various graph neural networks in the presence of features from various modalities. MM-Graph aims to foster research on multimodal graph learning and drive the development of more advanced and robust graph learning algorithms. By providing a diverse set of datasets and benchmarks, MM-Graph enables researchers to evaluate and compare their models in realistic settings, ultimately leading to improved performance on real-world applications that rely on multimodal graph data.
@@ 1 Introduction
Graphs are ubiquitous data structures that can effectively represent complex relationships and interactions between entities in various domains, such as social networks, biological systems, and recommendation systems. In real-world scenarios, these entities often possess rich semantic information in the form of unstructured data, such as images and text descriptions. Associating these unstructured data with the structured graph information is essential for tasks that require relevance search and information retrieval, such as recommendation systems [27, 3, 10].
Despite the importance of multimodal graph learning, existing graph learning benchmarks typically focus on investigating graphs with various connectivity patterns, such as long-range links, out-of-distribution scenarios, etc [5, 9, 19, 20, 13]. These benchmarks aim to evaluate the performance of graph neural networks (GNNs) and other graph learning algorithms on different graph structures and topologies. However, these benchmarks under-utilize the rich semantic information naturally present in each node, limiting their ability to evaluate graph learning algorithms in realistic settings.
With the prevalence of language modeling, there has been an outbreak in the effective integration of text and graph topology on text-attributed graphs, leading to a comprehensive benchmark on text-attributed graphs [33, 35, 15, 23]. While textual information undoubtedly plays a crucial role in understanding the data and making predictions for user-centric tasks, it is essential to recognize that other modalities, particularly visual information, can be helpful and provide complementary information that text alone cannot capture. For example, as shown in Figure  1(a), visually similar products may have distinct text features and ignoring such semantic gap between modalities may prevent GNNs from fully exploiting the relationships between multimodal semantics and structures.
Figure 1: Visualization of our Multimodal Graph Benchmark. All nodes of our benchmark have both visual and text features. (a) Amazon-Sports: The image and text come from the original image and title of the sports equipment. (b) Goodreads-LP: The image comes from the cover of the book. We do not show the text features of Goodreads-LP since the book description is very long. (c) Ele-fashion: The image and text come from the original image and title of the fashion product.
To address this limitation, we introduce the Multimodal Graph Benchmark (MM-Graph), the first comprehensive graph benchmark that incorporates both textual and visual information. MM-Graph consists of five graph learning datasets of various scales, designed for diverse downstream tasks such as node classification and link prediction. These datasets contain multimodal node features, including text and images, alongside the traditional graph structure. By providing a diverse set of datasets and benchmarks, MM-Graph enables researchers to evaluate and compare their models in realistic settings, fostering research on multimodal graph learning. MM-Graph also standardizes GNN architectures, feature encoders, and dataloaders and evaluators in order to enable a comprehensive evaluation on multimodal graph data
To show the importance of these datasets in for advancing multimodal graph learning , we adapt two state-of-the-art multimodal graph neural networks from the recommendation systems to the graph learning domain. These models have demonstrated strong performance in leveraging multimodal information for recommendation tasks. By applying them to the datasets in MM-Graph, we aim to provide a comprehensive evaluation of their effectiveness in various graph learning scenarios. We conduct an extensive study on the performance of MMGCN [31], MGAT [28], and other popular graph neural network models, such as GraphSAGE [11], in the presence of multimodal node features.
Surprisingly, we find that GNNs specifically designed for multimodal scenarios, such as MMGCN and MGAT, do not consistently outperform general-purpose GNN models like GraphSAGE across all datasets in MM-Graph. This finding highlights the need for a standard multimodal graph benchmark to rigorously evaluate and compare the performance of different models. MM-Graph fills this gap by providing a diverse set of datasets with varying scales, tasks, and modality combinations. By establishing a common ground for evaluation, MM-Graph enables researchers to gain a deeper understanding of the strengths and limitations of existing multimodal graph learning algorithms and identify areas for improvement. Moreover, MM-Graph opens up new research opportunities in multimodal graph learning. The availability of datasets with rich textual and visual information allows researchers to explore novel architectures and techniques that can effectively capture and integrate the semantic information from multiple modalities. This can lead to the development of more advanced and robust graph learning algorithms that can handle the complexities of real-world multimodal graph data.
Contributions and novelty. In this work, we introduce a multimodal graph benchmark known as MM-Graph. As design principles, we strive to create non-trivial, realistic data that comes from real-world applications and will benefit real-world deployments. Specifically, MM-Graph contains three link prediction datasets and two node classification datasets, as shown in Table  1. We summarize our novel contributions as follows: (1) To the best of our knowledge, no existing standardized graph benchmark provides both text and image features. This is important as images provide complementary context and details. While one can extract keywords and intent from text, images supply visual cues about objects, scenes, emotions, etc. that give richer understanding, especially for more abstract or complex tasks. (2) We provide a comprehensive analysis on the performance of various GNNs on MM-Graph, highlighting the need of developing new multimodal GNNs. (3) In terms of graph benchmarks, MM-Graph is the first benchmark to include not only various GNNs but also different ways of encoding node features, which presents an important and practical question of which feature encoder gives the best features for structure learning tasks.
Reproducibility: MM-Graph code, datasets and details are publicly available. The datasets are hosted on Huggingface. The evaluation code and baselines are hosted on GitHub with documentations.
@@
Table 1: Overview of the datasets in the proposed MM-Graph. We present five multimodal graph datasets with varying scale and tasks. All datasets are splitted in random. LP = Link Prediction. NC = Node Classification.
@@
Name Text Features Visual Features Proposed Task Metrics Scale Split Ratio
@@ 2 Related Work
Multimodal Graph Learning. While most existing multimodal learning approaches focus on 1-to-1 mappings between modalities, real-world data often exhibits more complex many-to-many relationships that can be represented as graphs. Multimodal graph learning (MMGL) aims to address this challenge by leveraging graph structures to learn from multimodal data with complicated relations among multiple multimodal neighbors  [6, 34]. Recent efforts in MMGL have focused on integrating GNNs with pretrained language models (LMs) [33]. Our work builds upon the recent advances in multimodal feature learning and MMGL by introducing a comprehensive benchmark that incorporates both structured and unstructured modalities. We adapt up-to-date multimodal feature learning frameworks and state-of-the-art multimodal GNNs and provide an extensive study on their performance in the presence of features from various modalities.
Benchmarks for graph representation learning. Several established graph benchmarks have been developed and widely adopted [20, 13, 7]. However, when it comes to learning on graphs with rich features, these benchmarks exhibit notable deficiencies. Firstly, a majority of these datasets suffer from the absence of raw feature information, limiting the investigation of attribute modeling’s effectiveness. Secondly, these datasets often neglect to explore the impact of feature modeling on GNNs. Thus, there is a compelling necessity to construct a comprehensive graph benchmark with rich multimodal features. Yan et al. proposed the first standardized graph benchmark with rich textual information, fostering the development of encompassing GNNs with prevalent language models [33]. However, we argue that it is critical to investigate modalities other than text, especially images in structure learning tasks and proposing the first standarized multimodal graph benchmark to foster the development of research in this direction.
3 MM-Graph: The Multimodal Graph Benchmark
In this section, we discuss the proposed 5 real-world datasets in MM-Graph. The basic information is shown in Table  1. We first provide diverse graph-related statistics so as to give a comprehensive understanding of the structural properties of the datasets. This information can guide the selection of appropriate graph learning algorithms and help interpret the performance of different models on each dataset.
@@ 3.1 Data Curation
We first introduce the data curation process of all datasets, as long as their license.
Amazon-Sports Amazon-Sports is a link prediction dataset, based on the Amazon-Review dataset [21, 12], where each node corresponds to a product on Amazon in the sports category and the link captures whether two products are co-purchased together. The text features are the titles of the products and the visual features are the high-resolution raw images of the products. In Figure  1(a), we show an example of nodes and edges in Amazon-Sports. The visual features are images of tents of various shapes under diverse backgrounds and text features are the title of the tents. Specifically, we use the data pre-processing script provided by Patton [16].
Amazon-Cloth Similar to Amazon-Sports, Amazon-Cloth is constructed according to the co-purchasing information of Amazon products in the clothes category. The text features are titles of products, e.g. Nike mens Revolution 6 Road Running and the visual features are corresponding images. We follow the same data curation procedure.
Goodreads-LP is based on the Goodreads Book Graph dataset  [29, 30]. Here we construct the graph in the way where each node corresponds to a book on Goodreads and the link captures if a user who likes this book will like the other book, following  [35]. The text features are the descriptions of the books and the visual features are the cover of the book. Nodes without images are removed.
Goodreads-NC is a node classification dataset, based on the Goodreads dataset [29, 30], where each node corresponds to a book on Goodreads and the link captures if a user who likes this book will like the other book. The text features are the descriptions of the books and the visual features are the cover image of the book. Nodes without images are removed.
Ele-fashion Ele-fashion is a node classification dataset, based on the Amazon-Fashion dataset [21, 12], where each node corresponds to a product and the link captures if a user who buy this product will also buy the other product. The text features are the titles of the products and the visual features are the images of the products.
Data Availability and Ethics. Our benchmark is organized from existing open source data, with proper open source licenses. Amazon-Sports and Amazon-Cloth are available with Apache License  [^1] https://github.com/PeterGriffinJin/Patton. Goodreads-LP, Goodreads-NC and Ele-fashion are released under MIT License  [^2] https://mengtingwan.github.io/data/goodreads.html. These do not involve interaction with humans or private data.
Table 2: Detailed graph-based statistics of each dataset proposed in MM-Graph. CC = Cluster Coefficient. RA = Resource Allocation. N/A = Not Applicable (Nodes do not have node labels). Detailed definition for each statistics provided in the Appendix.
@@ 3.2 Tasks
Link Prediction. The task is to predict new association edges given the training edges. The evaluation is based on how well a model ranks positive test edges over negative test edges. Specifically, for Amazon-Sports and Amazon-Cloth, we generate hard valid/test negatives using HeaRT  [18]. HeaRT is recognized as a better way of generating negatives for link prediction tasks in the community. For Goodreads-LP, since it takes more than 120 hours to generate hard negatives using HeaRT, we perform random sampling. Each positive edge in the validation/test set is ranked against 150 hard negative edges. For evaluation metrics, we report MRR, Hits@10, and Hits@1 - the three most commonly-used evaluation metrics for link prediction [13, 18]. For Amazon-Sports and Amazon-Cloth, edges are randomly split into train/valid/test splits according to 8/1/1 ratio. For Goodreads-LP, edges are randomly split according to 6/1/3 ratio. Validation and test edges are explicitly removed from the graphs to avoid any potential leakage [36].
Node Classification. The task is to predict the book category from $N$ available categories. We report accuracy following  [13, 33]. For Goodreads-NC, there are 10 categories such as History, Children and comics. For Ele-Fashion, there are 12 categories such as shoes, jewelry and dresses. For evaluation metrics, we report accuracy, as it’s the most common metric for node classification[13]. Nodes are randomly split into train/valid/test following 6/1/3 ratio.
@@ 3.3 Statistics
We provide graph-related statistics are presented in Table  2. The detailed descriptions for each metric is provided in the Appendix. These statistics provide valuable insights into the structural properties of the graphs in each dataset.
Amazon-Sports.There are 50,250 nodes with a total of 356,202 edges in Amazon-Sports, which is a small-scale link prediction dataset. The small scale makes Amazon-Sports great for trying it at first, as it is computation and memory efficient for most applications. The average degree is 14, which indicates that each node have 14 neighbors on average.
Amazon-Cloth.There are 125,839 nodes with a total of 951,271 edges in Amazon-Cloth. Amazon-Cloth constitutes a medium-scale link prediction dataset. Since the scale of this dataset is much larger than Amazon-Sports, Amazon-Cloth serves as a solid benchmark for real-world e-commerce based applications. The average degree is 15, which similar to Amazon-Sports. But the average clustering coefficient is much smaller. This indicates that nodes in Amazon-Cloth are more likely to scatter around, instead of forming communities despite its high average degree.
Goodreads-LP.There are 636,502 nodes with a total of 3,437,017 edges, which constitutes a large-scale link prediction dataset. The average degree is 10.
Goodreads-NC. There are 685,294 nodes with a total of 7,235,084 edges, which constitutes a large-scale node classification dataset. The average degree is 21, highest among all datasets.
Ele-fashion. There are 97,766 nodes with a total of 199,602 edges, which constitutes a medium-scale node classification dataset. The average degree is 4, which is the most sparse dataset among MM-Graph.

### v2

# Mosaic of Modalities: A Comprehensive Benchmark for Multimodal Graph Learning

@@ Abstract
Graph machine learning has made significant strides in recent years, yet the integration of visual information with graph structure and its potential for improving performance in downstream tasks remains an underexplored area. To address this critical gap, we introduce the Multimodal Graph Benchmark (MM-Graph), a pioneering benchmark that incorporates both visual and textual information into graph learning tasks. MM-Graph extends beyond existing text-attributed graph benchmarks, offering a more comprehensive evaluation framework for multimodal graph learning. Our benchmark comprises seven diverse datasets of varying scales (ranging from thousands to millions of edges), designed to assess algorithms across different tasks in real-world scenarios. These datasets feature rich multimodal node attributes, including visual data, which enables a more holistic evaluation of various graph learning frameworks in complex, multimodal environments. To support advancements in this emerging field, we provide an extensive empirical study on various graph learning frameworks when presented with features from multiple modalities, particularly emphasizing the impact of visual information. This study offers valuable insights into the challenges and opportunities of integrating visual data into graph learning.
@@ 1 Introduction
Graphs are powerful data structures that capture complex relationships across various domains, from social networks to recommendation systems. In real-world applications, these graph entities often possess rich, multimodal semantic information, including crucial visual data. The integration of visual information with structured graph data is particularly vital for tasks such as relevance search and information retrieval in recommendation systems [38, 4, 12].
Existing graph learning benchmarks have made significant strides in evaluating graph neural networks (GNNs) on various connectivity patterns and topologies [6, 11, 24, 27, 16], they often overlook the critical role of visual information. Moreover, recent advancements in language modeling have led to effective integration of textual information with graph structures, resulting in comprehensive benchmarks for text-attributed graphs [47, 50, 18, 30]. However, visual information offers unique and invaluable insights that text and structure alone cannot capture [44, 45]. Visual data can reveal similarities and relationships in artist styles that may not be apparent from textual or structural information alone. For instance, as illustrated in LABEL:fig:teaser, visually similar products may have distinct textual descriptions, creating a semantic gap that text-only GNNs might fail to bridge. By incorporating visual information, graph learning algorithms can better understand and exploit the intricate relationships between multimodal semantics and graph structures. This visual context can significantly enhance the performance of GNNs. Integrating visual data into graph learning frameworks enables a more holistic understanding of the underlying relationships in complex, real-world networks. It allows GNNs to capture nuanced information that may be missed when relying solely on textual or structural data, leading to more accurate and contextually relevant predictions.
To address the critical need for integrating visual data in graph machine learning, we introduce the Multimodal Graph Benchmark (MM-Graph), the first comprehensive graph benchmark that explicitly incorporates visual information alongside textual data. MM-Graph comprises seven diverse graph learning datasets of varying scales, specifically designed to evaluate two fundamental categories of graph tasks: node and link property predictions. These tasks require models to make predictions at different levels of graph granularity, leveraging the rich multimodal information available. A key feature of MM-Graph is its inclusion of high-resolution visual node features, which sets it apart from existing benchmarks. Each dataset within MM-Graph contains not only textual information and traditional graph structures but also associated images for nodes. This visual data provides crucial additional context that can significantly enhance the performance of graph learning algorithms in real-world scenarios.
While multimodal knowledge graphs (MMKGs) incorporating visual information exist, the community faces challenges with limited datasets, often with outdated or inaccessible image URLs and poor dataset quality [26, 36, 25, 3]. Recognizing the critical need for high-quality MMKG datasets that effectively integrate visual data [25], our MMKGs are sourced from CoDEx, a recent comprehensive knowledge graph benchmark known for its diverse and interpretable content and challenging link prediction tasks [36]. By enhancing CoDEx with curated visual information, our MMKGs offer higher quality datasets that address a significant gap in the MMKG community, providing a robust foundation for evaluating how graph learning models incorporate and benefit from visual features in knowledge graph scenarios. We highlight the key differences between MM-Graph and existing benchmarks in  Tab. 1.
Table 1: Qualitative comparison of existing benchmarks. Abbreviations: LP = Link Prediction. NC = Node Classification. KGC = Knowledge Graph Completion.
Benchmarks Language Vision NC LP KGC
GL-Bench [23]
CS-TAG [47]
VTKG [20]
CoDEx [35]
OGB [16]
MM-Graph
Contributions and novelty. In this work, we introduce a multimodal graph benchmark known as MM-Graph. MM-Graph fills this gap by providing a diverse set of datasets with varying scales, tasks, and modality combinations. By establishing a common ground for evaluation, MM-Graph enables researchers to reflect on the strengths and limitations of existing multimodal graph learning algorithms and identify areas for improvement. As design principles, we strive to create non-trivial, realistic data that comes from real-world applications and can benefit real-world deployments. Specifically, MM-Graph contains three link prediction datasets, two node classification datasets, and two knowledge graph completion datasets as shown in Tab. 2. We summarize our novel contributions as follows:
(1) Text- and Image- Attributed Graph(TIAG) Benchmarks. First standardized benchmark providing both textual and visual features across various graph learning tasks, enabling a more holistic understanding of complex data relationships.
(2) Benchmarking Multimodal GNNs. Comprehensive performance analysis of diverse methods on MM-Graph, highlighting the need for more advanced multimodal Graph Neural Networks (GNNs).
(3) Exploring Multimodal Feature Encodings. Pioneering investigation of various node feature encoding strategies (e.g. ImageBind), addressing the challenge of optimal feature representation in graph learning tasks.
@@
Table 2: Overview of the datasets in our proposed MM-Graph benchmark. We present five multimodal graph datasets with varying scale and tasks. All datasets are split at random. Abbreviations: LP = Link Prediction. NC = Node Classification. KGC = Knowledge Graph Completion. OR = Original.
@@
Name Text Features Visual Features Task Metrics Scale Split Ratio
@@
MM-CoDEx-s Entity Description Entity Images KGC MRR, Hits@K Small OR
MM-CoDEx-m Entity Description Entity Images KGC MRR, Hits@K Medium OR
@@ 2 Related Work
Multimodal Graph Learning. While most existing multimodal learning approaches focus on 1-to-1 mappings between modalities, real-world data often exhibits more complex many-to-many relationships that can be represented as graphs. Multimodal graph learning (MMGL) aims to address this challenge by leveraging graph structure to learn from multimodal data with complicated relations among multiple multimodal neighbors  [7, 48, 8, 14]. Recent efforts in MMGL have focused on integrating GNNs/Knowledge Graph Embeddings (KGEs) with pretrained language models (LMs) [47]. Our work builds upon the recent advances in multimodal feature learning and MMGL by introducing a standardized benchmark that incorporates both structured and unstructured modalities.
Benchmarks for graph representation learning. Several established graph benchmarks have been developed and widely adopted [27, 16, 9]. However, when it comes to learning on graphs with rich features, these benchmarks exhibit notable deficiencies. Firstly, these datasets suffer from the absence of raw features, limiting the investigation of attribute modeling. Secondly, these datasets often neglect to explore the impact of feature modeling. Thus, there is a compelling necessity to construct a comprehensive graph benchmark with rich natural features. Yan et al. proposed a graph benchmark with rich textual information, fostering the development of integrating GNNs with prevalent language models [47]. However, we argue that it is critical to investigate modalities other than text, especially images in structure learning tasks and proposing the first standarized multimodal graph benchmark to foster the development of research in this direction. The key differences between MM-Graph and existing benchmarks are highlighted in  Tab. 1.
3 MM-Graph: Multimodal Graph Benchmark
In this section, we present the proposed 7 real-world datasets in MM-Graph. The general information about them is given in  Tab. 2. We first introduce the data curation process of all datasets, as well as their license. For all datasets described below, we use Beautiful Soup to crawl images associated with the nodes in the graphs [33].
@@ 3.1 Data Curation
Amazon-Sports/Amazon-Cloth are link prediction datasets, based on the Amazon-Review dataset [28, 15], where each node corresponds to a product on Amazon in the sports category and the link captures whether two products are co-purchased together. The text features are the titles of the products and the visual features are the high-resolution raw images of the products. In LABEL:fig:teaser(a), we show an example of nodes and edges in Amazon-Sports. The visual features are images of tents of various shapes under diverse backgrounds and text features are the titles of the tents. We preprocess the data by extracting the meta information of each item as well as their co-purchasing information.
Goodreads-LP/Goodreads-NC are based on the Goodreads Book Graph dataset  [41, 42]. Here we construct the graph such that each node corresponds to a book on Goodreads and the link captures if a user who likes this book will like the other book, following  [50]. The text features are the descriptions of the books and the visual features are the the book covers. Nodes without images are removed.
Ele-fashion is a node classification dataset, based on the Amazon-Fashion dataset [28, 15], where each node corresponds to a product and a link captures that the two products are copurchased. The text features are the titles of the products and the visual features are the images of the products.
MM-CoDEx-s/MM-CoDEx-m is a knowledge graph completion dataset, based on the CoDEx-s dataset [36]. The text features come from the description of the Wikipedia article. For visual features, we crawl the corresponding images of the entity (e.g., person/place/item) from Wikipedia.
Data Availability and Ethics. Our benchmark is organized from existing open source data, with proper open source licenses. Amazon-Sports and Amazon-Cloth and Ele-Fashion are available with Apache License  [^1] https://github.com/PeterGriffinJin/Patton. Goodreads-LP, Goodreads-NC, MM-CODEX-S and MM-CODEX-M are released under MIT License  [^2] https://mengtingwan.github.io/data/goodreads.html, https://github.com/tsafavi/codex/tree/master. These do not involve interaction with humans or private data.
Table 3: Detailed graph-based statistics of datasets proposed in MM-Graph. CC = Cluster Coefficient. RA = Resource Allocation. N/A = Not Applicable (Nodes do not have class labels). Detailed definition for each graph statistics or property are provided in the Appendix.
@@
Table 4: Detailed MMKG statistics of datasets proposed in MM-Graph. (+): Positive (true) triples. (-): Verified negative (false) triples.
Name Entities Relations Train (+) Valid (+) Test (+) Valid (-) Test (-)
MM-CoDEx-s 1,383 39 14,298 784 802 1028 1074
MM-CoDEx-m 7,697 51 47,617 2,628 2,595 4,721 4,746
@@ 3.2 Tasks
Link Prediction. The task is to predict new association edges given the training edges. The evaluation is based on how well a model ranks positive test edges over negative test edges. Specifically, for Amazon-Sports and Amazon-Cloth, we generate hard valid/test negatives using HeaRT [22]. HeaRT is recognized as a better way of generating negatives for link prediction in the community. For Goodreads-LP, since it takes more than 120 hours to generate hard negatives using HeaRT, we perform random sampling. Each positive edge in the validation/test set is ranked against 150 hard negative edges. For evaluation metrics, we report MRR, Hits@10, and Hits@1, the three most commonly-used evaluation metrics for link prediction [16, 22]. For Amazon-Sports and Amazon-Cloth, edges are randomly split into train/valid/test splits according to 8/1/1 ratio. For Goodreads-LP, edges are randomly split according to 6/1/3 ratio. Validation and test edges are explicitly removed from the graphs to avoid any potential leakage [51].
Node Classification. For Goodreads-NC, the task is to predict the category of items from 10 available categories, such as History, Children and comics. For Ele-Fashion, there are 12 categories such as shoes, jewelry and dresses. For evaluation metrics, we report accuracy, as it’s the most common metric for node classification[16, 47]. Nodes are randomly split into train/valid/test following 6/1/3 ratio.
Knowledge Graph Completion. The task is to predict missing links in a knowledge graph given entities and relations. Similar to link prediction, the evaluation is based on how well a model ranks positive test edges over negative test edges. Specifically, we use the splits and negatives designed in  [36] to ensure the scope and level of difficulty of the task. Entities without multimodal information (i.e., entities that do not have descriptions or images) are filtered out. For evaluation metrics, we report MRR, Hits@10, Hits@3 and Hits@1.
@@ 3.3 Statistics
We present detailed statistics for each dataset in Tab. 3 and Tab. 4, and additional descriptions of each metric and property are provided in the Appendix. These statistics provide valuable insights into the structural properties of the graphs.
Our multimodal graph benchmark encompasses datasets spanning various scales and characteristics. The datasets range from small-scale (MM-CoDEx-s) to large-scale (Goodreads-NC) with varying entity and relation counts. The diversity in graph sizes, densities, and average degrees provides a comprehensive testbed for evaluating multimodal graph learning approaches across different computational complexities.